# Dynamic Content Moderation in Livestreams: Combining Supervised Classification with MLLM-Boosted Similarity Matching

### v1

@@ Abstract.
$c$s: General and reference General conference proceedings
@@ 1. Introduction
In this work, we propose a hybrid moderation framework that integrates two complementary paradigms: preset violation detection based on supervised multiclass classification and reference-based similarity matching based on video-clip retrieval and re-ranking. Our architecture is designed to address the dual need for precision in known violation categories and flexibility in identifying emerging or ambiguous edge cases. The system ingests multimodal live-stream inputs, processing them through a dual-path architecture. The first path employs a supervised classification model trained on labeled examples to detect explicit, high-confidence violations. The second path leverages a similarity matching engine that compares incoming content to a curated set of known policy violations using semantic and perceptual embeddings, enabling the system to generalize to previously unseen behaviors through nearest-neighbor retrieval. While our offline benchmarks show that a finetuned multimodal large language–vision model (MLLM) would offer the strongest performance, it is not feasible to deploy such a model in production due to resource constraints. Instead, we adopt a knowledge distillation approach, using the MLLM as a teacher to guide and enhance the performance of preset violation detection model and the re-ranking model during training.
@@
  • Production-Scale Deployment with Comprehensive Impact Evaluation: We deploy our system on a large-scale live-streaming platform processing millions of hours of content. Our empirical evaluation demonstrates substantial gains in coverage at hight precision, and large scale online A/B experiments show a reduction of unwanted livestream views by $6\%$ to $8\%$, demonstrating measurable impact in real-world production.
@@ 2.2. Multimodal Learning for Video Understanding
Recent works have applied large language models (LLMs) and have shown they can be effective at content moderation for language-related tasks (Palla et al., 2025; Zeng et al., 2024; Kumar et al., 2024). However, for online platforms, the challenge arises from the complexity other media, which involve complex multi-modal signals including visual, auditory, and speech. Multi-modal LLMs (MLLMs) have explored techniques to integrate additional modalities, such as video or audio, from prevailing proprietary models such as GPT4o (Team, 2024b) and Gemini (Team, 2024a), to recent open-source models such as InternVL (Chen et al., 2024), QwenAudio (Chu et al., 2023), and Llava-One-Vision (Li et al., 2024b). As such, MLLMs have began to be applied in the context of content moderation, by integrating image/video information (Lu et al., 2025b; AlDahoul et al., 2024), as well as audio (Luo et al., 2024). The key challenge lies in effective integration of multi-modal information while leveraging existing knowledge from the pretraining of these large models.
@@ 2.3. Leveraging LLMs for Knowledge Distillation
Figure 1. Overview of our proposed hybrid moderation framework. Incoming live-stream clips are segmented into 20-second windows and processed into multimodal inputs (frames, audio, text). The system follows two parallel paths: (1) Preset Violation Detection using a supervised multiclass classification model trained on labeled data, and (2) Reference Matching through a similarity-based retrieval system. Retrieved candidates are refined by a multimodal re-ranking model that evaluates semantic alignment across modalities. Enforcement decisions for Live Streams are made predefined decision rules.
@@ 3.3. Reference-based Similarity Matching Pipeline
To further enhance learning, we apply knowledge distillation (Figrue 5 in Appendix) using a fine-tuned large language-vision model, the LLaVa-One-Vision(Li et al., 2024a), which provides richer semantic supervision. The student model (multimodal re-ranking model) is trained to align with the teacher’s outputs, enabling it to better capture nuanced cross-modal relationships.
@@ 4.7. Online Performance
[…]
@@ Appendix A Appendix
Clip-Match Aggregation Algorithm
@@
Knowledge Distillation of Re-ranking model

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783936isbn: 979-8-4007-2258-5/2026/08submissionid: v1ads0109ccs: General and reference General conference proceedings
@@ 1. Introduction
In this work, we propose a hybrid moderation framework that integrates two complementary paradigms: preset violation detection based on supervised multiclass classification and reference-based similarity matching based on video-clip retrieval and re-ranking. Our architecture is designed to address the dual need for precision in known violation categories and flexibility in identifying emerging or ambiguous edge cases. The system ingests multimodal live-stream inputs, processing them through a dual-path architecture. The first path employs a supervised classification model trained on labeled examples to detect explicit, high-confidence violations. The second path leverages a similarity matching engine that compares incoming content to a curated set of known policy violations using semantic and perceptual embeddings, enabling the system to generalize to previously unseen behaviors through nearest-neighbor retrieval. The two paths are complimentary as there is a large portion of detections are exclusive. Specifically, reference matching contributes approximately 22% additional coverage beyond the classification branch. While our offline benchmarks show that a finetuned multimodal large language–vision model (MLLM) would offer the strongest performance, it is not feasible to deploy such a model in production due to resource constraints. Instead, we adopt a knowledge distillation approach, using the MLLM as a teacher to guide and enhance the performance of preset violation detection model and the re-ranking model during training.
@@
  • Production-Scale Deployment with Comprehensive Impact Evaluation: We deploy our system on a large-scale live-streaming platform processing millions of hours of content. Our empirical evaluation demonstrates substantial gains in coverage at high precision, and large scale online A/B experiments show a reduction of unwanted livestream views by $6\%$ to $8\%$, demonstrating measurable impact in real-world production.
@@ 2.2. Multimodal Learning for Video Understanding
Recent works have applied large language models (LLMs) and have shown they can be effective at content moderation for language-related tasks (Palla et al., 2025; Zeng et al., 2024; Kumar et al., 2024; Zhu et al., 2025). However, for online platforms, the challenge arises from the complexity of the media, which involve complex multi-modal signals including visual, auditory, and speech. Multi-modal LLMs (MLLMs) have explored techniques to integrate additional modalities, such as video or audio, from prevailing proprietary models such as GPT4o (Team, 2024b) and Gemini (Team, 2024a), to recent open-source models such as InternVL (Chen et al., 2024), QwenAudio (Chu et al., 2023), and Llava-One-Vision (Li et al., 2024b). As such, MLLMs have began to be applied in the context of content moderation, by integrating image/video information (Lu et al., 2025b; AlDahoul et al., 2024), as well as audio (Luo et al., 2024). The key challenge lies in effectively utilizing the MLLM to real-world use cases in a resource-efficient way.
@@ 2.3. Leveraging LLMs for Knowledge Distillation
Figure 1. Overview of our proposed hybrid moderation framework. Incoming live-stream clips are segmented into 20-second windows and processed into multimodal inputs (frames, audio, text). The system follows two parallel paths: (1) Preset Violation Detection using a supervised multiclass classification model trained on labeled data, and (2) Reference Matching through a similarity-based retrieval system. Retrieved candidates are refined by a multimodal re-ranking model that evaluates semantic alignment across modalities. Enforcement decisions for Live Streams are made of predefined decision rules.
@@ 3.3. Reference-based Similarity Matching Pipeline
To further enhance learning, we apply knowledge distillation (Figrue 6 in Appendix) using a fine-tuned large language-vision model, the LLaVa-One-Vision(Li et al., 2024a), which provides richer semantic supervision. The student model (multimodal re-ranking model) is trained to align with the teacher’s outputs, enabling it to better capture nuanced cross-modal relationships.
@@ 4.7. Online Performance
The online A/B experiment was conducted over a few weeks with $10\%$ traffic assigned to each of the experiment groups. The evaluation was performed using a two-sample z-test for proportions to measure the change in view rate of violated live streams.
  • Metric: $1.2\%$ decrease in violation-related user views.
  • Standard deviation: $0.055$ over the sampled population.
  • Confidence interval ($95\%$): $0.1476\%,-0.087\%]$.
  • P-value: $\sim 0$.
  • Minimum Detectable Effect (MDE): $0.04\%$.
@@
Hit Latency:
Percentile (ms)
Module Average (ms) 50th 90th 99th
Preset Violation 417.55 350.06 703.42 1000.00
Ref. Matching 3980.00 4120.00 5490.00 6310.00
Table 8. Runtime latency evaluation per 20 second clip for the two pipelines. On average, the preset violation branch’s latency is about 0.4 s, while the reference-matching branch’s latency is about 4 s, which enables very quick detection for live streams.
To monitor the impact and effectiveness of our pipeline, a long-term A/B backtest experiment is carried out to monitor the performances of the preset violation detection pipeline and reference matching pipeline. The preset violation detection pipeline alone leads to about $4\%$ reduction in user views of unwanted live streams. Introducing reference matching pipeline further reduces the user views of unwanted live streams down by about $2\%$ to $4\%$, demonstrating how the enhanced reference matching pipeline could effectively complement the preset violation detection pipeline to improve overall moderation efficacy.
To minimize false positives and over-moderation, the system employs several safeguards:
  • Model thresholds are tuned at P90 precision to minimize overkill in automated moderation.
  • Human moderators review detections in lower precision bands (P70-P85) for verification before enforcement.
  • For affected creators, an appeal mechanism routes disputed cases directly to the human moderation team, where decisions are re-evaluated based on both model outputs and contextual review.
@@ Appendix A Appendix
A.1. Clip-Match Aggregation Algorithm
@@
A.2. More Ablation Results
Aggregation Clip Length
Metric No Aggr. 5 10 15
Model Hit Rate 0.39% 0.48% 0.41% 0.36%
True Detection Rate 0.35% 0.44% 0.38% 0.34%
Overkill Rate 0.04% 0.04% 0.03% 0.01%
Table 9. Sensitivity of the system to hyperparameters (e.g., clip length, embedding dimension). Aggregation improves recall (0.35% to 0.44%) at similar overkill rates, allowing us to lower the similarity threshold and achieve a higher true-positive rate at similar overkill levels, while longer windows balance precision and recall by slightly reducing both true-positives and false-positives.
Aggr. AP F1 P70 P75 P80 P85 P90
No aggr. 75.84% 73.61% 75.63% 71.40% 66.70% 62.93% 47.37%
1 min 76.84% 74.51% 76.85% 71.44% 68.25% 67.69% 46.85%
2 min 77.68% 75.01% 77.79% 72.84% 68.96% 66.15% 46.11%
3 min 78.15% 75.42% 79.18% 73.20% 70.15% 69.81% 45.61%
4 min 78.30% 75.83% 78.84% 73.54% 70.55% 70.30% 45.91%
5 min 78.49% 76.06% 79.62% 73.83% 70.01% 70.55% 46.10%
Table 10. Score aggregation of the reference matching branch: sliding windows of different durations. Aggregation boosts AP(75.8 to 78.5) and F1(73.6 to 76.1), showing temporal smoothing improves stability. In deployment, we adopt a simple sliding-window aggregation with mean pooling.
Model AP F1
Gemini-2.5-Pro 8.29% 15.32%
Our Approach 75.84% 73.61%
Table 11. Comparison with a pretrained MLLM (Gemini-2.5-Pro) on the classification branch. The pretrained model exhibits extremely low precision and F1 when applied directly to the moderation domain, likely due to hallucination tendencies and a lack of task-specific grounding.
Model AP F1
Gemini-2.5-Pro 30.28% 35.84%
Our Approach 74.82% 73.49%
Table 12. Comparison with a pretrained MLLM (Gemini-2.5-Pro) on the reference-matching branch. While the pretrained model performs better than in the classification setting, it still significantly underperforms our approach, highlighting the importance of task-specific modeling and grounding for reliable moderation.
A.3. Exploration of Score Distribution
(a) Preset Violation Detection model
(b) Reference Matching model
Figure 5. We observe that raising precision from $80\%$ to $90\%$ requires a steep threshold increase, as ’violation’ and ’non-violation’ scores heavily overlap in this range, indicating most hard cases lie here. And, both precision and recall remain stable around $75-80\%$, with minor fluctuations, demonstrating model stability.
A.4. Knowledge Distillation of Re-ranking model